OpenStudy (anonymous):

Whoever answers this will get a medal. (: It takes a florist 3 hrs 15 min to make 3 small centerpieces and 3 large centerpieces. It takes 6 hrs 20 mins to make 4 small centerpieces. How long does it take to make each centerpiece and each large centerpiece?

OpenStudy (anonymous):

So whats the answer?

OpenStudy (anonymous):

65

OpenStudy (pixiedust1):

Oh wait! Is your question saying that it made both in that time, or just one size?

OpenStudy (anonymous):

Sorry, I'm very bad at understanding algebra. I don't know how to solve the large centerpiece. Lol im sorry and i think in that time

OpenStudy (pixiedust1):

Okay, and I read the question wrong... so i messed up too! i realized that they wanted BOTH in that amount of time! So hold on while I rework the question.

OpenStudy (anonymous):

Alright!!

OpenStudy (pixiedust1):

let the time taken to make large piece be x hrs let the time taken to make the small piece by y hrs 3x + 3y = 3 1/4 = 13/4 (#1) 7x + 4y = 6 1/3 = 19/3 (#2) #1 times 4: 12x+12y = 13 #2 times 3: 21x+12y=19 subtract them ... 9x = 6 x = 2/3 hour or 40 minutes in #1 3(2/3) + 3y = 13/4 2 + 3y = 13/4 3y = 13/4 - 2 = 5/4 y = 5/12 hours or 25 minutes

OpenStudy (pixiedust1):

Okay, sorry about that

OpenStudy (pixiedust1):

I believe that's all right. Is that one of your answer options?

OpenStudy (anonymous):

whats the answer? Its not a multiple choice question. :( and dont worry about it, its no problem!

OpenStudy (pixiedust1):

So the large take 40 minutes, and the small take 25 minutes

OpenStudy (anonymous):

thank you so much!

OpenStudy (pixiedust1):

No problem! Have an awesome day! :D